Circular No. 6510

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

CYGNUS X-1 S. N. Zhang and C. R. Robinson, Universities Space Research Associ-

ation; and W. Cui, Massachusetts Institute of Technology (MIT), report: “We have detected the 5.6-day orbital period of the blackhole candidate Cyg X-1 in the ASM/RXTE data, confirming a similar detection with the Ginga ASM in its low-energy band during the x-ray hard state (Kitamoto et al., submitted to Ap.J.). The modulation is only detectable in the hard state (before ∼ 1996 May 17 and after ∼ Aug. 10) and is strongest in the lowest ASM energy band (1.3–3.0 keV), with a full amplitude of ∼ 10 percent and x-ray flux minimum near superior conjunction of the compact object. This seems to indicate that the modulation is due to an asymmetry in the photoelectric absorption. However, the lack of any significant orbital modulation in the soft state suggests that another effect may be responsible for the detected modulation, unless the power-law component is much more heavily absorbed than the blackbody component, since the ASM 1.3–3.0keV detector counting rate is dominated by the power-law component in the hard state and by the blackbody component in the soft state.”

SUPERNOVAE G. V. Williams, Smithsonian Astrophysical Observatory, reports his

measures of the positions of seven supernovae discovered by J. Mueller. Given below are the supernova designation, position (internal uncertainty < 0′′.4) as measured from the Digital Sky Survey (based on original Palomar Sky Survey), offset from host galaxy’s center, and reference to the discovery report. The nucleus of the host galaxy (UGC 465) for SN 1954ac is illdefined.

SN  α_{2000}  δ_{2000}  Offset  IAUC
1950O 16^{h}15^{m}09.^{s}39 + 18^{o}57′14′′.4 10′′.0 E, 12′′.8 S 5536
1953J  0 44 55.66 + 16 56 10.8 14′′.5 E, 4′′.5 S 5140
1953L 5 15 46.52 + 71 27 41.6 3′′.6 W, 8′′.6 N 5466
1954ac 0 43 50.78 + 32 50 32.9 0′′.5 W, 38′′.2 S 5140
1954ad 8 32 46.37 + 0 13 33.7 7′′.2 W, 5′′.9 S 5175
1955R 12 34 42.49 + 49 19 54.8 3′′.8 E, 26′′.0 N 5525
1955S 15 36 40.61 + 43 32 38.0 16′′.2 W, 15′′.8 N 5761
1996 November 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green